# FusionBench: A Comprehensive Benchmark of Deep Model Fusion

### v3

@@ Abstract
Deep model fusion is an emerging technique that unifies the predictions or parameters of several deep neural networks into a single model in a cost-effective and data-efficient manner. This enables the unified model to take advantage of the original models’ strengths, potentially exceeding their performance. Although a variety of deep model fusion techniques have been introduced, their evaluations tend to be inconsistent and often inadequate to validate their effectiveness and robustness against distribution shifts. To address this issue, we introduce FusionBench, which is the first comprehensive benchmark dedicated to deep model fusion. FusionBench covers a wide range of tasks, including open-vocabulary image classification, text classification, and text-to-text generation. Each category includes up to eight tasks with corresponding task-specific models, featuring both full fine-tuning and LoRA fine-tuning, as well as models of different sizes, to ensure fair and balanced comparisons of various multi-task model fusion techniques across different tasks, model scales, and fine-tuning strategies. We implement and evaluate a broad spectrum of deep model fusion techniques. These techniques range from model ensemble methods, which combine the predictions to improve the overall performance, to model merging, which integrates different models into a single one, and model mixing methods, which upscale or recombine the components of the original models. FusionBench now contains 26 distinct tasks, 74 fine-tuned models, and 16 fusion techniques, and we are committed to consistently expanding the benchmark with more tasks, models, and fusion techniques. In addition, we offer a well-documented set of resources and guidelines to aid researchers in understanding and replicating the benchmark results. This includes detailed documentation, code examples, and tutorials, making FusionBench a user-friendly and accessible platform for both beginners and experienced researchers. The FusionBench homepage is available at https://github.com/tanganke/fusion_bench.
@@ 1 Introduction
In recent years, a new paradigm called “learn from model” has emerged in the field of deep learning, which focuses on leveraging the knowledge embedded in existing models to develop new ones [67]. This paradigm has been widely adopted in various scenarios, such as model tuning [18, 8], model distillation [21], model pruning [16, 3], model editing [38, 66], and so on. Among these methods, deep model fusion is particularly appealing. It merges the parameters or predictions of multiple models to create a more robust and efficient unified model. Due to its effectiveness and scalability, many new techniques for deep model fusion have recently been proposed [34].
@@
Deep model fusion offers both scalability and data efficiency by utilizing the knowledge embedded in pre-existing models, rather than requiring training from scratch. This approach significantly accelerates model development, making it a practical solution in the current era dominated by large foundation models. Despite its potential, the evaluation of deep model fusion techniques often suffers from inconsistency and inadequacy. Standardized assessments are lacking, making it challenging to verify their effectiveness and robustness. The potential reasons for this inconsistency include the rapid development of new techniques, the absence of standardized tasks and models, and the variety of settings (such as different fine-tuning strategies). Additionally, challenges in implementing or replicating prior work contribute to these inconsistencies.
To tackle these challenges, we develop the first comprehensive benchmark dedicated to deep model fusion, called FusionBench. FusionBench is built to be a modular and extensible platform, comprising three core modules: the Algorithm Module, the Model Pool Module, and the Task Pool Module. Each module is configurable, allowing users to easily customize and manage the benchmark settings. In summary, our main contributions are four-fold:

### v4

# FusionBench: A Unified Library and Comprehensive Benchmark for Deep Model Fusion

@@ Abstract
Deep model fusion is an emerging technique that unifies the predictions or parameters of several deep neural networks into a single better-performing model in a cost-effective and data-efficient manner. Although a variety of deep model fusion techniques have been introduced, their evaluations tend to be inconsistent and often inadequate to validate their effectiveness and robustness. We present FusionBench, the first benchmark and a unified library designed specifically for deep model fusion. Our benchmark consists of multiple tasks, each with different settings of models and datasets. This variety allows us to compare fusion methods across different scenarios and model scales. Additionally, FusionBench serves as a unified library for easy implementation and testing of new fusion techniques. FusionBench is open source and actively maintained, with community contributions encouraged.
[Uncaptioned image] Code Repository: https://github.com/tanganke/fusion_bench
[Uncaptioned image] Online Documentation: https://tanganke.github.io/fusion_bench
[Uncaptioned image] HuggingFace Collections (models & datasets): https://huggingface.co/tanganke
Keywords: Deep Model Fusion, Model Merging, Multi-Task Learning, Benchmark
@@ 1 Introduction
In recent years, a new paradigm called “learn from model” has emerged in the field of deep learning, which focuses on leveraging the knowledge embedded in existing models to develop new ones (zheng2023learn). This paradigm has been widely adopted in various scenarios, such as model tuning (he2022masked; chung2024scaling), model distillation (hinton2015distilling), model pruning (han2015deep; asif2020ensemble), model editing (mitchell2021fast; zhang2024comprehensive), and so on. Among these methods, deep model fusion is particularly appealing. It merges the parameters or predictions of multiple models to create a more robust and efficient unified model. Due to its effectiveness and scalability, many new techniques for deep model fusion have recently been proposed (li2023deep).
@@
Deep model fusion offers both scalability and data efficiency by utilizing the knowledge embedded in pre-existing models, rather than requiring training from scratch. This approach significantly accelerates model development, making it a practical solution in the current era dominated by large foundation models. Despite its potential, the evaluation of deep model fusion techniques often suffers from inconsistency and inadequacy. Current libraries like MergeKit (goddard2024arcee) are primarily focused only on large language model merging and do not provide evaluation frameworks. Standardized assessments across different tasks and models are lacking, making it challenging to evaluate the effectiveness and robustness of deep model fusion techniques. The potential reasons for this inconsistency include the rapid development of new techniques, the absence of standardized tasks and models, and the variety of settings (such as different fine-tuning strategies). Additionally, challenges in implementing or replicating prior work contribute to these inconsistencies. A more detailed comparison with existing model fusion methods and toolkits is provided in Section A.
To tackle these challenges, we develop the first comprehensive benchmark dedicated to deep model fusion, called FusionBench. FusionBench is designed as a modular and flexible framework with three primary components: the Algorithm Module, the Model Pool Module, and the Task Pool Module as shown in Figure 1. Each component offers full configurability, enabling straightforward integration of novel algorithms, models, and evaluation tasks. Through Hydra-based configuration management, all experimental setups are streamlined via a single unified interface. Furthermore, FusionBench provides extensive documentation, hands-on tutorials, illustrative code samples, and ready-to-use model collections.
2 FusionBench
FusionBench is available for installation through two simple methods: using pip with ‘pip install fusion-bench’ or cloning the repository directly with ‘git clone’. The general framework of the modularized FusionBench codebase is shown in Figure 1, which consists of three primary elements: Algorithm Module, Model Pool Module, and Task Pool Module. In Section 2.1, we introduce the three major components, which is designed to be flexible and modular, allowing users to easily run experiments and evaluate the performance of model fusion algorithms. In Section 2.2, we introduce the implemented model fusion algorithms. Finally, in Section 2.3, we discuss the documentation and tutorials provided to help users understand the benchmark and effectively use the codebase. In Appendix C, we provide a flowchart to illustrate the process of running experiments and evaluating the merged models.
Main Program Pipeline.
The typical pipeline consists of the following steps: (1) Users configure the experiment by selecting a fusion algorithm, models to merge, and optional evaluation tasks. (2) The system loads the specified models, algorithm, and evaluation tasks. (3) The algorithm merges the selected models, with an option to save the resulting model. (4) When evaluation tasks are specified, the merged model is tested and results are reported.
2.1 Components
As shown in Figure 1, the codebase is composed of three primary components: Algorithm Module, Model Pool Module, and Task Pool Module, which are responsible for implementing the model fusion algorithms, managing the models and datasets used during the algorithm execution, and managing the tasks and evaluation metrics to be evaluated, respectively. Additionally, we provide a unified command line interface (CLI) ‘fusion_bench’ to facilitate the use of the codebase and to enable users to easily run experiments.
  • Algorithm Module is the core component, which contains the implementation of all model fusion algorithms. Each algorithm is implemented as a separate Python class inheriting from the base class fusion_bench.BaseAlgorithm. The algorithm classes are flexible - they can be used through our CLI or in custom code.
  • Model Pool Module provides a unified interface for managing models and datasets.
  • Task Pool Module evaluates model performance on different tasks, where each task combines a dataset with specific evaluation metrics.
2.2 Implemented Algorithms
Table 1: Implemented algorithms in FusionBench.
Method Requirement
Model Ensemble Methods
Simple Ensemble (sagi2018ensemble) -
Weighted Ensemble (sagi2018ensemble) hyperparameter search
Max-Model Predictor (wu2019heterogeneous) -
Model Merging Methods
Simple Average / Modelsoups (wortsman2022model) -
Weighted Average (matena2022merging) hyperparameter search
Fisher Merging (matena2022merging) compute weights on labeled data
RegMean (jin2022dataless) compute weights on labeled data
RegMean++ (nguyen2025regmean++) compute weights on labeled data
Concrete Subspace (tang2023concrete) test-time adaptation training
Task Arithmetic (ilharco2022editing) hyperparameter search
Ties-Merging (yadav2023resolving) hyperparameter search
Task-Wise AdaMerging (yang2023adamerging) test-time adaptation training
Layer-Wise AdaMerging (yang2023adamerging) test-time adaptation training
Representation Surgery (yang2024representation) unlabeled data samples
TALL mask (wang2024localizing) -
Task Singular Vectors (gargiulo2024task) -
Isotropic Merging (marczak2025no) -
OPCM (tang2025merging) -
FW-Merging (chen2025fw) labeled data samples
RanDeS (zhou2025randes) -
Model Mixing Methods
MoE-based Upscaling (komatsuzaki2022sparse) pre-training to recover performance
MoE-based Merging (komatsuzaki2022sparse) training on the combined model
Depth Upscaling (kim2023solar) pre-training to recover performance
Model Recombination (hu2023fedmr) training on the combined model
Weight-Ensemble MoE (tang2024merging) test-time adaptation training, vision tasks
WE-MoE V2 (shen2024efficient) test-time adaptation training, vision tasks
Pareto-Driven Merging (tang2024towards) training datasets
SMILE Upscaling (tang2024smile) -
Model Compression Methods
BitDelta (liu2024bitdelta) calibration data
Magnitude Pruning (frankle2020linear) -
Wanda Pruning (sun2023simple) calibration data
SparseGPT Pruning (frantar2023sparsegpt) calibration data
Expert Pruning (lu2024not) calibration data
Our benchmark incorporates a comprehensive collection of cutting-edge model fusion algorithms, spanning three key categories: ensemble methods, merging methods, and mixing methods. Additionally, we have recently extended FusionBench to include model compression methods, which, while not strictly fusion techniques, are complementary approaches that can be applied to reduce the computational overhead of models or compress individual models before and after fusion. Additionally, we evaluated their prominence in the research literature and their practical utility. The complete listing of these implemented algorithms can be found in Table 1 and FusionBench is under active development and will be updated to include the latest and most effective model fusion algorithms.
2.3 Online Documentation and Tutorials
We provide detailed documentation and tutorials on our project website to help users get started. These resources cover the basics of model fusion, how to run experiments, and how to evaluate results. We also include example pre-run results showing how different fusion algorithms perform across various tasks.
3 Conclusion and Discussion
FusionBench offers a robust framework for evaluating deep model fusion algorithms and a unified library for developing new ones. Its scalable and extensible architecture simplifies the creation of deep model fusion techniques. Furthermore, FusionBench provides curated datasets and models to ensure fair comparisons. With comprehensive documentation and tutorials, it is accessible to both beginners and experienced researchers. We encourage the community to use this benchmark to advance the field of deep model fusion.
Acknowledgments and Disclosure of Funding
This work is supported by the National Natural Science Foundation of China (Grant No. 62225113, U23A20318, U2336211, 62576364 and 62276195), the Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China (Grant No. JYB2025XDXM704), the Shenzhen Basic Research Project (Natural Science Foundation) Basic Research Key Project (NO. JCYJ20241202124430041), the Foundation for Innovative Research Groups of Hubei Province (Grant No. 2024AFA017) and the Science and Technology Major Project of Hubei Province (Grant No. 2024BAB046). Dr. Tao’s research is partially supported by NTU RSR and Start Up Grants. The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.
The appendix is organized into several sections, each providing additional insights and details related to different aspects of the main work.
Appendix A Related Work of Deep Model Fusion
Figure 2: A taxonomy of deep model fusion techniques.
Deep model fusion, often referred to in the literature simply as “model fusion”, aims to combine the outputs or parameters of multiple deep neural networks into a unified model with enhanced performance, efficiency, or robustness. Different researchers may categorize these techniques in various ways based on their understanding and points of view. Here, we propose a taxonomy that divides these techniques into three major categories: Model Ensemble, Model Merging, and Model Mixing. Each of these categories approaches model fusion from a unique perspective, offering distinct advantages and applicability. In the following, we provide detailed explanations, formal definitions, and analyze their strengths and weaknesses. A visualization of the taxonomy is shown in Figure 2.
A.1 A Taxonomy of Deep Model Fusion Techniques
Model Ensemble methods combine the predictions of multiple models to improve the overall performance of a machine learning system (sagi2018ensemble), where the collective knowledge is often more accurate and reliable than that of any individual model. Mathematically, given a set of $N$ models $\{f_{1},f_{2},\dots,f_{N}\}$, which can be homogeneous or heterogeneous, we use their predictions to obtain a global prediction $y=\mathcal{A}_{ensemble}(x;f_{1},f_{2},\dots,f_{N};w)$, where $\mathcal{A}_{ensemble}$ is an ensemble algorithm and $w$ are the algorithmic parameters. Each model $f_{i}$ can also be associated with a specification to indicate its weight or importance in the ensemble (pathak2010multiparty; zhouLearnwareFutureMachine2016; wu2021model; tangImprovingHeterogeneousModel2023a). Ensemble methods are widely used and effective in improving performance but are often expensive to use and manage. Recent research has also investigated efficient techniques for model ensembles (wen2020batchensemble; chen2023split; allingham2021sparse).
Model Merging methods integrate the parameters of multiple models into a unified model, enhancing efficiency in terms of inference cost and storage, and enabling scalable model fusion. Given a set of $N$ isomorphic models $\{f_{i}(\cdot;\theta_{i})\}_{i=1}^{N}$, each parameterized with $\theta_{i}$, we merge them into a single model with parameters $\theta=\mathcal{A}_{merging}(\theta_{1},\theta_{2},\dots,\theta_{N};w)$, where $\mathcal{A}_{merging}:\mathbb{R}^{N\times d}\rightarrow\mathbb{R}^{d}$ is a merging algorithm and $w$ are the algorithmic parameters. The merged model can be expressed as $f(\cdot;\theta)$. This method can be implemented through linear interpolation in parameter space (wortsman2022model; ilharco2022editing; yadav2023resolving; matena2022merging; yu2024language; chronopoulou2023adaptersoup; rame2024rewarded; ortiz2024task; liu2023tangent), leveraging mode connectivity (draxler2018essentially; frankle2020linear; benton2021loss; garipov2018loss; qu2024rethink), aligning features, parameters or gradients (liu2022deep; ainsworthGitReBasinMerging2023; jin2022dataless; tam2024merging; stoica2023zipit; jang2023personalized; daheim2023model; yang2024representation), subspace-based methods (tang2023concrete; wang2024localizing; yi2024safety; zhu2024model; xu2024weight), and ensemble distillation (wan2024knowledge; wan2024fusechat). Model merging methods are often performed in a data-efficient manner, the algorithmic parameters $w$ can also be learned during test time via test-time adaptation (TTA) training or meta-learning for a more seamless merging (yang2023adamerging; tang2023concrete).
Model Mixing methods fuse the components of multiple models to create a new heterogeneous model, which can be more flexible and adaptive than the original models. Mathematically, given a set of $N$ models $\{f_{i}(\cdot;\theta_{i})\}_{i=1}^{N}$, each parameterized with $\theta_{i}\in\mathbb{R}^{d}$, we mix their components to obtain a new model with parameters $\Theta=\mathcal{A}_{mixing}(\theta_{1},\theta_{2},\dots,\theta_{N};w)\in\mathbb{R}^{d^{\prime}}$, where $\mathcal{A}_{mixing}:\mathbb{R}^{N\times d}\mapsto\mathbb{R}^{d^{\prime}}$ is a mixing algorithm and $w$ is the algorithmic parameters. The mixed model can be expressed as $F(\cdot;\Theta)$, which often has more parameters than the original models, and thus can be more expressive and powerful to capture the underlying patterns in the data. Model mixing methods can be implemented through layer recombinations (hu2023fedmr; jiang2024evomerge), model stitching (lenc2015understanding; moschella2022relative), or upscale to create a Mixture of Experts (MoE)-based sparse model (komatsuzaki2022sparse; ye2023taskexpert; tang2024merging; lu2024twin; dai2024deepseekmoe; zhao2024loraretriever; ostapenko2024towards; tang2024smile; yadav2024survey).
A.2 Related Benchmarks and Toolkits
Although several model fusion methods have been proposed, comprehensive benchmarks and unified toolkits are still lacking in this field. A recent notable work, MergeKit (goddard2024arcee), provides a collection of model fusion techniques specifically designed for merging large language models (LLMs), focusing primarily on methods for Transformer-based architectures. While MergeKit excels in providing practical tools for LLM merging, its scope is inherently limited to a single domain (natural language processing) and a specific model family (Transformers). Similarly, MergeBench (he2025mergebench) is targeted exclusively at evaluating LLM merging techniques, providing benchmarks only within the language modeling domain.
In contrast, FusionBench addresses several critical gaps in the current landscape, making FusionBench a more generalized and versatile platform for systematically assessing the performance of model fusion approaches across various domains and architectures: (1) Broader Domain Coverage: Unlike existing tools that focus solely on NLP, FusionBench spans both computer vision and natural language processing domains, enabling cross-domain evaluation and comparison. (2) Comprehensive Taxonomy: While MergeKit and MergeBench concentrate primarily on model merging methods, FusionBench encompasses a complete taxonomy including model ensemble, model merging, and model mixing methods, providing a unified framework for all major fusion paradigms. (3) Research-Oriented Evaluation: FusionBench is designed as a research platform that includes diverse fine-tuned models, standardized evaluation protocols, and comparative analysis utilities across different architectures (CNNs, Vision Transformers, CLIP models, etc.) and tasks. (4) Algorithmic Diversity: Beyond basic merging techniques, FusionBench incorporates advanced methods such as subspace-based fusion, gradient alignment, and mixture-of-experts approaches.
Appendix B Project Structure
To provide a clear overview of FusionBench’s organization, we present its comprehensive project structure in Listing 1. This structure reflects the modular design philosophy of our benchmark, with distinct directories for configuration management, documentation, example implementations, core functionality, and testing.
1|-- config/ # Configuration files managed by Hydra
2| |-- method/ # Configs for implemented algorithms
3| |-- modelpool/ # Configs for model pool
4| |-- taskpool/ # Configs for task pool
5| |-- model/ # Configs for models
6| |-- dataset/ # Configs for datasets
7|-- docs/ # Documentation (mkdocs)
8|-- examples/ # Example scripts and notebooks
9| |-- {method_name}/ # Method-specific examples
10|-- fusion_bench/ # Main package
11| |-- method/ # Fusion algorithm implementations
12| | |-- {method_name}/ # Method-specific code
13| |-- modelpool/ # Model pool implementations
14| |-- taskpool/ # Task pool implementations
15| |-- ... # Models, datasets, and other utilities
16|-- tests/ # Unit tests
Listing 1: Project structure of FusionBench.
Figure 3: Flowchart of FusionBench.
Appendix C FusionBench Architecture and Implementation Guide
Figure 3 illustrates the steps to fuse a model in FusionBench across three main stages: creating a model pool, instantiating and running an algorithm, and (optional) model evaluation.
Pseudocode Example.
The pseudocode for the pipeline is shown in Algorithm 1.
Algorithm 1 Pseudocode for FusionBench main pipeline
1:Load and parse configuration from YAML files
2:Instantiate fusion_algorithm, model_pool, and (optionally) task_pool
[…]
4:Save merged_model if save path is specified
5:if task pool is specified then
6:  $\textit{report}\leftarrow\textit{task\_pool.evaluate}(\textit{merged\_model})$
7:  Save report if save path is specified
8:end if
This design allows users to flexibly experiment with different algorithms, models, and tasks by simply modifying the configuration, without changing the core program logic.
C.1 Implementing Customized Algorithm in FusionBench
Here, we provide an example of implementing a customized algorithm in Listing 2. The derived algorithm class should inherit from the “BaseAlgorithm” class and implement the “run” method. The “run” method takes a “BaseModelPool” object as input, which is responsible for managing the models and dataset to be loaded. The “run” method should be implemented to fuse the models and return the fused model.
1from fusion_bench import BaseAlgorithm, BaseModelPool
2
3class DerivedAlgorithm(BaseAlgorithm):
4 """
5 An example of a derived algorithm class.
6 """
7
8 # _config_mapping maps the attribution to the corresponding key in the configuration file.
9 # this is optional and can be used to serialize the object to a configuration file.
10 # ‘self.config.hyperparam_1‘ will be mapped to the attribute ‘hyperparam_attr_1‘.
11 _config_mapping = BaseModelFusionAlgorithm._config_mapping | {
12 "hyperparam_attr_1": "hyperparam_1",
13 "hyperparam_attr_2": "hyperparam_2",
14 }
15
16 def __init__(self, hyperparam_1, hyperparam_2, **kwargs):
17 super().__init__(**kwargs)
18 self.hyperparam_attr_1 = hyperparam_1
19 self.hyperparam_attr_2 = hyperparam_2
20
21 def run(self, modelpool: BaseModelPool):
22 # modelpool is an object that responsible for managing the models and dataset to be loaded.
23 # implement the fusion algorithm here.
24 raise NotImplementedError(
25 "DerivedAlgorithm.run() is not implemented."
26 )
Listing 2: Implementing new algorithm in FusionBench.
As an illustration, we can implement a deep model fusion algorithm that merges models through parameter averaging (wortsman2022model). The following code snippet demonstrates this approach (Listing 3):
1class SimpleAverageAlgorithm(
2 BaseAlgorithm,
3 SimpleProfilerMixin,
4):
5 @torch.no_grad()
6 def run(self, modelpool: Union[BaseModelPool, Dict[str, $n$.Module]]):
7 """
8 Fuse the models in the given model pool using simple averaging.
9
10 This method iterates over the names of the models in the model pool, loads each model, and appends it to a list.
11 It then returns the simple average of the models in the list.
12
13 Args:
14 modelpool: The pool of models to fuse.
15
16 Returns:
17 The fused model obtained by simple averaging.
18 """
19 if isinstance(modelpool, dict):
20 modelpool = BaseModelPool(modelpool)
21
22 log.info(
23 f"Fusing models using simple average on {len(modelpool.model_names)} models."
24 f"models: {modelpool.model_names}"
25 )
26 sd: Optional[StateDictType] = None
27 forward_model = None
28 merged_model_names = []
29
30 for model_name in modelpool.model_names:
31 with self.profile("load model"):
32 model = modelpool.load_model(model_name)
33 merged_model_names.append(model_name)
34 print(f"load model of type: {type(model).__name__}")
35 with self.profile("merge weights"):
36 if sd is None:
37 # Initialize the state dictionary with the first model’s state dictionary
38 sd = model.state_dict(keep_vars=True)
39 forward_model = model
40 else:
41 # Add the current model’s state dictionary to the accumulated state dictionary
42 sd = state_dict_add(sd, model.state_dict(keep_vars=True))
43 with self.profile("merge weights"):
44 # Divide the accumulated state dictionary by the number of models to get the average
45 sd = state_dict_div(sd, len(modelpool.model_names))
46
47 forward_model.load_state_dict(sd)
48 # print profile report and log the merged models
49 self.print_profile_summary()
50 log.info(f"merged {len(merged_model_names)} models:")
51 for model_name in merged_model_names:
52 log.info(f" - {model_name}")
53 return forward_model
Listing 3: Implementation of simple average algorithm in FusionBench.
C.2 Customizing Model Pools in FusionBench
The Model Pool Module serves as an abstraction layer that manages model loading, saving, and dataset handling during fusion operations. To extend FusionBench to new model architectures, users need to implement a custom model pool class that inherits from BaseModelPool and defines architecture-specific operations. This design principle ensures that fusion algorithms remain architecture-agnostic and can be applied universally across different model types without modification.
Extensibility to Modern Architectures. FusionBench’s modular architecture facilitates straightforward integration of emerging model architectures. The key advantage of this design is that once a model pool class is implemented for a particular architecture, all existing fusion algorithms in the benchmark become immediately applicable to that architecture. This extensibility has enabled us to expand beyond our initial focus on established models (CLIP-ViT, ResNet-50, GPT-2, and Flan-T5) to encompass state-of-the-art architectures across multiple domains.
For computer vision tasks, we have integrated modern architectures including ConvNeXt (liu2022convnet) and DINOv2 (oquab2024dinov), both of which represent significant advances in efficient visual representation learning. These models can be loaded through dedicated model pool classes (ConvNeXtForImageClassificationPool and DINOv2ForImageCla-
ssificationPool
) with comprehensive online documentation available on our project website.
For natural language processing tasks, our CausalLMPool class provides broad compatibility with contemporary large language model architectures. The CausalLMPool leverages the HuggingFace Transformers library, allowing seamless integration of new architectures as they are released.
Appendix D Fine-Tuned Model Collections in FusionBench
This section outlines the comprehensive model collections provided in FusionBench, including their experimental configurations and performance evaluation results across various single-task specializations.